ACCURATE MOLECULAR CONSTANTS, POTENTIAL CURVE AND BORN-OPPENHEIMER BREAK-DOWN CORRECTION FUNCTIONS FOR X^{1}Σ_{g}^{+} MgH and MgD

\underline{GANG LI}, JENNING Y. SETO, PETER F. BERNATH and ROBERT J. LE ROY, Guelph-Waterloo Centre for Graduate Work in Chemistry and Biochemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada; RAM S. RAM, Department of Chemistry, University of Arizona, Tucson, AZ 85721, USA.

New high resolution Fourier Transform spectra of the 0–3, 0–4, 1–3, 1–4 and 1–5 bands of the B′ ^{2}Σ^{+} − X^{2}Σ^{+} transition of MgH, together with lines from sunspot spectra coupling B′–state levels v′ = 0 & 1 to X–state levels v′′ = 3 − 8 , combined with older B′ − X band data^{a} involving v′′(X) = 3 − 9 and v′(B′) = 0 − 9 and ground-state 2–1 and 1–0 infrared data, yield a description of the ground state for v′′ = 0 − 2 and v′′ = 3 − 9 . The v′′ = 2 − 3 gap was bridged using transitions from the 1–2 and 1–3 bands of the A^{2}Π − X^{2}Σ^{+} spectrum. In order to avoid complications due to perturbations in the excited state, all of these electronic band data were re-arranged and treated as fluorescence series into the ground state. The resulting data set consisted of a total of 4140 transitions for six isotopomers of MgH and MgD. We have performed two types of combined isotopomer analyses of these data. (i) A fit to empirical Dunham-type expansions which included hydrogenic Born-Oppenheimer breakdown correction terms required 54 expansion parameters, plus the 701 “fluorescence series” origins. However, the resulting empirical centrifugal distortion constants will be unreliable for extrapolation to J values significantly higher than those included in the data set. (ii) An equally good fit to an analytic model potential energy function plus adiabatic and non-adiabatic radial Born-Oppenheimer breakdown correction functions required only 21 potential and correction-function parameters, plus the 701 “fluorescence series” origins. The resulting potential function has the MLJ form,^{b} and was constrained to have the correct C_{6}/R^{6} long-range behaviour, so it should prove highly reliable for extrapolations in v or J. The parameter-fit and potential-fit computer programs used for this analysis were DSParFit and DSPotFit.^{c}

a W.J. Balfour and H.M. Cartwright, Can.J.Phys. 54, 1898 (1976); W.J. Balfour and B. Lindgren, Can.J.Phys. 56, 767 (1978).

b P.H. Hagigeorgiou and R.J. Le Roy, J. Chem. Phys. 112, 3949 (2000).

c Program source code and manuals available through the “Computer Programs” link on the www site http://leroy.uwaterloo.ca.